# Language Repository for Long Video Understanding

### v1

@@ 4.1 Writing to repository
In each subsequent iteration $k$, previous repo entries of iteration $k-1$ are re-combined into longer chunks and processed in the same way, generating information for higher semantic-levels.
@@
Here, re-chunk($\cdot$) denotes the creation of longer chunks within the repository. More specifically, we simply concatenate (denoted by []delimited-[][$\cdot$][ ⋅ ]) all entries from previous iteration, and split them again into fewer number of chunks (hence, longer chunk size). Note that $i^{\prime}$ in the above equation is not the same as the previous chunk indexing $i$, as we may have different (usually, fewer) number of chunks in each subsequent iteration.
Each write operation involves two stages: (1) Grouping redundant text, and (2) Rephrasing, which are detailed below.
@@ 4.1.0.1 Grouping redundant text:
Given textual descriptions of a video chunk (i.e., captions in the first write iteration, or previous repo descriptions in subsequent iterations), we plan to identify most-similar ones and merge them as a single description. Without loss of generality, let us consider the first write iteration, for which the input is in the form of $c_{i}=\{c_{ij}\;|\;j=1,\cdots,p\}$. Inspired by [5], we first split the captions of each chunk into two sets, namely, source (src) captions $c_{\text{src},i}$ and destination (dst) captions $c_{\text{dst},i}$. Let us drop the chunk index $i$ for brevity. Here, dst captions $c_{\text{dst}}$ are sampled uniformly distributed across the temporal span of a chunk, while all the rest are considered as src captions $c_{\text{src}}$ (see Fig. 2).
@@
Here, we usually have fewer dst captions (i.e., $|c_{\text{dst}}|<|c_{\text{src}}|$). Next, we embed all captions using a text-encoder (e.g. CLIP [59]), and compute the cosine similarity of each pair between src-dst sets to find most-similar matches.
$\text{sim}_{\text{src-dst}}$ $=\texttt{sim}(\texttt{encode}(c_{\text{src}}),\;\texttt{encode}(c_{\text{dst}})).$ (4)
@@
$c_{\text{grp}}$ $=\texttt{group}(c_{\text{dst}},\;c_{\text{src}},\;\text{sim}_{\text{src-dst}},\;x).$ (5)
Here $x\%$ is a pre-defined hyperparameter (e.g. 50%). Finally, such grouped descriptions go through a rephrasing operation prior to entering the repository.
Figure 3: LLM prompt templates in LangRepo: Here, we show the zero-shot prompt templates used for rephrasing ($\texttt{template}_{\texttt{reph}}$) and summarizing ($\texttt{template}_{\texttt{sum}}$) operations. Rephrase prompt needs a list of grouped captions as input, while its output adheres to more-strict requirements (e.g. same order, same number of list items) needed for correct parsing. Summarize prompt takes in each repository entry and generates a more-flexible (i.e., open-ended) output, while also conditioning on the question.
@@ 4.1.0.2 Rephrasing:
Grouped captions $c_{\text{grp}}$ of each chunk are rephrased via an LLM-call. This allows redundant information within each group to be dropped, while generating a concise and coherent description. We first form a list of grouped captions, where each list item corresponds to a single group (i.e., a dst caption and any one or more src captions matched to it), and feed it to the LLM, wrapped in a rephrasing-template ($\texttt{template}_{\texttt{reph}}$) as shown in Fig. 3 (top).
${c_{\text{reph}}}$ $=\texttt{rephrase}(\texttt{template}_{\texttt{reph}}(c_{\text{grp}},\;l)).$ (6)
Here, $l$ denotes the number of grouped captions in the given chunk. The LLM output (${c_{\text{reph}}}$) is restricted to be a list in the same order with the same number of items, where each item is a single concise sentence. Finally, such rephrased descriptions together with other metadata such as timestamps ($t$) and number of occurrences ($o$) are written in the repository.
$r^{1}$ =$=\{(c_{\text{reph},j},\;t_{j},\;o_{j})\;|\;j=1,\;\cdots,\;p^{\prime}\}.$ (7)
Note that here $p^{\prime}<p$ as we have grouped and rephrased $x\%$ of most-similar captions. Alongside each description in a repository entry, $t$ maintains a growing list of timestamps corresponding to its founding captions, whereas the occurrences counter ($o$) keeps track of the number of such captions. A qualitative example of a repository entry is given in Fig. 4.
In subsequent iterations, the same operations apply when writing multi-scale entries. The only difference is the change in input, which now constitutes of previous repo entries re-combined into high-level chunks (i.e., $c\rightarrow\texttt{re-chunk}([\cdots,\;r^{k-1},\;\cdots])$ in Eq. 3). Each new iteration generates information corresponding to a higher semantic-level (i.e., going from short-range to long-range dependencies), forming our multi-scale language representation.
@@ 4.2 Reading from repository
As we make a single VQA prediction for a given long video— instead of making predictions every chunk— our read operation (read-from-repo) is applied only after fully-forming the multi-scale repository (i.e., after all write iterations). The repo entries after $K$ write iterations can be denoted as $\{r^{k}\;|\;k=1,\cdots,K\}$ where each scale ($r^{k}$) may consist of multiple entries $\{\cdots,\;r^{k}_{i-1},\;r^{k}_{i},\;r^{k}_{i+1},\;\cdots\}$. When reading, we generate summaries for each entry in the repo separately, allowing it to focus on varying temporal spans. More specifically, each entry goes through a summarizing-template ($\texttt{template}_{\texttt{sum}}$) as shown in Fig. 3 (bottom), and the resulting prompt is fed to the LLM.
@@
ans $=\texttt{vqa}([\cdots,d^{k}_{i},\cdots]).$ (9)
4.3 Design decisions
4.3.0.1 Similarity-based pruning:
4.3.0.2 Processing videos as chunks:
Our decision to consume longer videos as chunks is motivated by prior work [79, 63]. It allows us to not lose short-term details, while also keeping track of long-term dependencies via multi-scale processing. Additionally, although not explored in the scope of this paper, such a setup integrates well with temporally-fine-grained prediction tasks, where an LLM needs to make multiple predictions over time.
4.3.0.3 Choice of metadata:
4.3.0.4 All-textual repository:
Instead of being a latent representation [79, 63, 3], our LangRepo is all-textual. This promotes interpretability for human observers, while also being a more-natural form of structure for LLM-based processing. Additionally, our implementation can be formulated to be zero-shot, without requiring any training or finetuning.
4.3.0.5 Classifier for close-ended VQA:
The standard multiple-choice question-answering setup considers a generative classifier. Meaning, an LLM is prompted to generate the correct answer option among multiple-choices, directly as next-token prediction. Another approach used in NLP literature is the log-likelihood based classification. Here, the LLM is prompted separately for each of the multiple choices with a template such as “Question: Answer-option”. The choice that maximises the log-likelihood of predicted tokens (i.e., tokens corresponding to Answer-option) is selected as the correct answer. This is a more-natural setup for close-ended VQA since it avoids hallucination. Among these classifiers, we find the latter to be better-performing. Yet, it is more-sensitive to the prompt template. We direct the reader to supplementary A.1 for more details.
@@ 5 Experiments
5.1 Implementation details
In our experiments, we rely on frame or short-clip captions pre-extracted using VLLMs, as given in [92]. For EgoSchema [51], LaViLa [96] is used to generate captions at 1fps. For NExT-QA [80] data used in other benchmarks, LLaVA-1.5 [49] generates captions at 0.5fps. By default, we use either Mistral-7B [29] (w/ 7B parameters) or Mixtral-8$\times$7B [30] (w/ 12B active parameters) as our LLMs. As the text encoder in similarity-based pruning, we use CLIP-L/14 [59]. Note that all the models used in our framework are open-source and within a reasonable model-scale, promoting further research in academic settings. We do zero-shot inference on all datasets without any finetuning, evaluating the performance on long-term video VQA benchmarks.
5.2 Dataset details
@@ 5.2.0.1 EgoSchema:
EgoSchema [51] is a long-video VQA dataset derived from Ego4D [22], by semi-automatically generating and verifying mutliple-choice questions. Each video is 3 minutes long, and the questions are filtered to have a long temporal certificate (i.e., at least 100 seconds of a video is required to answer the question). The public validation subset consists of 500 videos, each with 5 answer-choices and the correct ground-truth answer. The fullset of 5K videos is hosted as a Kaggle challenge. There is no training split for this dataset, motivating zero-shot evaluation. In our experiments, we report the performance for both subset and fullset.
@@ 5.2.0.2 NExT-QA:
NExT-QA [80] is a popular VQA dataset with videos up to 2 minutes long, at an average of 44 seconds. It consists of 52k open-ended questions and 48k close-ended questions (i.e., multiple-choice with 5 answer options) in total. The questions are further classified into temporal, causal, or descriptive categories, to enable evaluation of different reasoning capabilities of models. We consider zero-shot evaluation on validation set of 5k multiple-choice questions, and report the performance on each category.

### v2

@@ 4.1 Writing to repository
In each subsequent iteration $k+1$, previous repo entries of iteration $k$ are re-combined into longer chunks and processed in the same way, generating information for higher semantic-levels.
@@
[…]
Here, re-chunk($\cdot$) denotes the creation of longer (and, fewer, i.e., $m<n$) chunks within the repository. More specifically, we simply concatenate (denoted by []delimited-[][$\cdot$][ ⋅ ]) all entries from previous iteration, and split them again into fewer number of chunks (hence, longer chunk size). Note that $i^{\prime}$ in the above equation is not the same as the previous chunk indexing $i$, as we may have different (usually, fewer) number of chunks in each subsequent iteration. Each write operation involves two stages: (1) Grouping redundant text, and (2) Rephrasing, which are detailed below.
@@ 4.1.0.1 Grouping redundant text:
Given textual descriptions of a video chunk (i.e., captions in the first write iteration, or previous repo descriptions in subsequent iterations), we plan to identify most-similar ones and merge them as a single description. Without loss of generality, let us consider the first write iteration, for which the input is in the form of $c^{0}_{i}=\{c^{0}_{ij}\;|\;j=1,\cdots,p\}$. Inspired by (Bolya et al., 2022), we first split the captions of each chunk into two sets, namely, source (src) captions $c^{0}_{\text{src},i}$ and destination (dst) captions $c^{0}_{\text{dst},i}$. Let us drop the chunk index ($i$) and iteration index (00) for brevity. Here, dst captions $c_{\text{dst}}$ are sampled uniformly distributed across the temporal span of a chunk, while all the rest are considered as src captions $c_{\text{src}}$ (see Fig. 2 top-left).
@@
Here, we usually have fewer dst captions (i.e., $|c_{\text{dst}}|<|c_{\text{src}}|$). Next, we embed all captions using a text-encoder (e.g. CLIP (Radford et al., 2021)), and compute the cosine similarity of each pair between src-dst sets to find most-similar matches.
$\text{sim}_{\text{src-dst}}$ $=\texttt{similarity}(\texttt{encode}(c_{\text{src}}),\;\texttt{encode}(c_{\text{dst}}))\;.$ (5)
@@
$c_{\text{grp}}$ $=\texttt{group}(c_{\text{dst}},\;c_{\text{src}},\;\text{sim}_{\text{src-dst}})\;.$ (6)
Here, an additional hyperparameter (i.e., $x$) decides the grouping ratio. Finally, such grouped descriptions go through a rephrasing operation prior to entering the repository.
Figure 3: LLM prompt templates in LangRepo: Here, we show the zero-shot prompt templates used for rephrasing ($\texttt{template}_{\texttt{reph}}$) and summarizing ($\texttt{template}_{\texttt{sum}}$) operations. Rephrase prompt needs a list of grouped captions as input, while its output adheres to more-strict requirements (e.g. same order, same number of list items) needed for correct parsing. Summarize prompt takes in each repository entry and generates a more-flexible (i.e., open-ended) output, while optionally conditioning on the question.
@@ 4.1.0.2 Rephrasing:
Grouped captions $c_{\text{grp}}$ of each chunk are rephrased via an LLM-call. This allows redundant information within each group to be dropped, while generating a concise and coherent description. We first form a list of grouped captions, where each list item corresponds to a single group (i.e., a dst caption and any one or more src captions matched to it), and feed it to the LLM, wrapped in a rephrasing-template ($\texttt{template}_{\texttt{reph}}$) as shown in Fig. 3 (top-left).
${c_{\text{reph}}}$ $=\texttt{rephrase}(\texttt{template}_{\texttt{reph}}(c_{\text{grp}}))\;.$ (7)
Here, the LLM output (${c_{\text{reph}}}$) is restricted to be a list in the same order with the same number of items, where each item is a single concise sentence. Finally, such rephrased descriptions together with other metadata such as timestamps ($t$) and number of occurrences ($o$) are written in the repository.
$r$ =$=\{(c_{\text{reph},j},\;t_{j},\;o_{j})\;|\;j=1,\;\cdots,\;p^{\prime}\}\;.$ (8)
Note that here $p^{\prime}<p$ as we have grouped and rephrased a pre-defined ratio (e.g. 50%) of most-similar captions. Alongside each description in a repository entry, $t$ maintains a list of timestamps corresponding to its founding captions, whereas the occurrences counter ($o$) keeps track of the number of captions grouped together. A qualitative example of a repository entry is given in Fig. 4.
In subsequent iterations, the same operations apply when writing multi-scale entries. The only difference is the change in input, which now constitutes of previous repo entries re-combined into high-level chunks (i.e., $c^{0}\rightarrow c^{k}$). Each new iteration generates information corresponding to a higher semantic-level (i.e., going from short-range to long-range dependencies), forming our multi-scale language representation.
@@ 4.2 Reading from repository
As we make a single VQA prediction for a given long video— instead of making predictions every chunk— our read operation (read-from-repo) is applied after fully-forming each scale of multi-scale repository (i.e., after writing all chunks). The repo entries from $K$ scales can be denoted as $\{r^{k}\;|\;k=0,\cdots,K\}$ where each scale ($r^{k}$) may consist of multiple entries $\{\cdots,\;r^{k}_{i-1},\;r^{k}_{i},\;r^{k}_{i+1},\;\cdots\}$. When reading, we generate summaries for each entry in the repo separately, allowing it to focus on varying temporal spans. More specifically, each entry goes through a summarizing-template ($\texttt{template}_{\texttt{sum}}$) as shown in Fig. 3 (bottom), and the resulting prompt is fed to the LLM.
@@
[…]
@@ 5 Experiments
In our experiments, we rely on captions pre-extracted using VLLMs, as given in (Zhang et al., 2023a). As for the LLM, we use either Mistral-7B (Jiang et al., 2023) (w/ 7B parameters) or Mixtral-8$\times$7B (Jiang et al., 2024) (w/ 12B active parameters) by default. As the text encoder in similarity-based pruning, we use CLIP-L/14 (Radford et al., 2021). Note that all the models used in our framework are open-source and within a reasonable model-scale, making our work accessible even in academic settings. We do zero-shot inference on all datasets without any finetuning, evaluating the performance on long-form video VQA benchmarks.
For evaluations, we consider four challenging long-video VQA benchmarks in our evaluations. EgoSchema (Mangalam et al., 2024) derived from Ego4D (Grauman et al., 2022), consists of 3-minute long clips, each with a question and 5 answer-choices. Its public validation subset consists of 500 videos, whereas the held-out fullset has 5K videos. NExT-QA (Xiao et al., 2021) contains videos up to 2 minutes long (at an average of 44 seconds), annotated with 52k open-ended questions and 48k close-ended questions (i.e., multiple-choice with 5 answer options). The questions are further classified into temporal, causal, or descriptive categories, to evaluate different reasoning capabilities of models. We consider zero-shot evaluation on the validation set. IntentQA (Li et al., 2023a) is based on the same NExT-QA videos, yet focuses more on intent-related questions (e.g. why?, how? or before/after) with a total of 16k multiple-choice questions on 4.3k videos. Here, we consider zero-shot setting on the test set. NExT-GQA (Xiao et al., 2023a) is a visually-grounded VQA dataset with 10.5K temporal grounding annotations, where we consider zero-shot inference similar to (Zhang et al., 2023a), on the test split.
5.1 Main results
Table 2: Results on EgoSchema (Mangalam et al., 2024): We present comparisons with state-of-the-art models on EgoSchema subset (500-videos) and fullset (5000-videos). We focus on the zero-shot setting. LangRepo shows a strong performance at its scale.
{tabu}
lcccc Model Video Pretrain Params  Subset (%)  Fullset (%)
finetuned
MC-ViT-L (Balažević et al., 2024) ✓ 424M 62.6 44.4
ImageViT (Papalampidi et al., 2023) ✓ 1B 40.8 30.9
ShortViViT (Papalampidi et al., 2023) ✓ 1B 47.9 31.0
LongViViT (Papalampidi et al., 2023) ✓ 1B 56.8 33.3
zero-shot (with proprietary LLMs)
Vamos (Wang et al., 2023) ✓ 175B - 41.2
Vamos (Wang et al., 2023) ✓ 1.8T - 48.3
LLoVi (Zhang et al., 2023a) ✗ 175B 57.6 50.3
ProViQ (Choudhury et al., 2023) ✗ 175B - 57.1
MoReVQA (Min et al., 2024) ✗ 340B - 51.7
LVNet (Park et al., 2024)$<$1.8T    68.2 61.1
VideoAgent (Wang et al., 2024b) ✗ 1.8T 60.2 54.1
VideoAgent (Fan et al., 2024) ✗ 1.8T 62.8 -
IG-VLM (Kim et al., 2024) ✗ 1.8T - 59.8
VideoTree (Wang et al., 2024e) ✗ 1.8T 66.2 61.1
LifelongMemory (Wang et al., 2024d) ✗ 1.8T 68.0 62.1
zero-shot (with open-source LLMs)
VIOLET(Fu et al., 2023) ✓ 198M - 19.9
InternVideo (Wang et al., 2022a) ✓ 478M - 32.1
FrozenBiLM (Yang et al., 2022) ✓ 890M - 26.9
SeViLA (Yu et al., 2024) ✓ 4B 25.7 22.7
Tarsier (Wang et al., 2024a) ✓ 7B 56.0 49.9
VideoChat2 (Li et al., 2024a) ✓ 7B 63.6 54.4
VideoLLaMA 2 (Cheng et al., 2024) ✓ 12B - 53.3
Vamos (Wang et al., 2023) ✓ 13B - 36.7
InternVideo2 (Wang et al., 2024c) ✓ 13B - 60.2
Tarsier (Wang et al., 2024a) ✓ 34B 68.6 61.7
mPLUG-Owl (Ye et al., 2023b) ✗ 7B - 31.1
Mistral (Jiang et al., 2023) ✗ 7B 48.8 -
LLoVi (Zhang et al., 2023a) ✗ 7B 50.8 33.5
LangRepo (ours) ✗ 7B 60.8 38.9
LangRepo (ours) ✗ 12B 66.2 41.2
@@ 5.1.0.1 EgoSchema:
In Table 5.1, we present the VQA performance of LangRepo on standard EgoSchema (Mangalam et al., 2024) splits, comparing with other state-of-the-art frameworks. Here, we focus on zero-shot evaluation, yet also report finetuned setting (i.e., any downstream-data-specific training) for completeness. We consider Mistral-7B (Jiang et al., 2023) and Mixtral-8$\times$7B (Jiang et al., 2024) as the choice of LLMs in our setup, both with reasonable model scales (7B and 12B active parameters, respectively). We de-emphasize the comparisons with models having significantly-higher #parameters (e.g. 175B GPT-3.5, or 1.8T GPT-4 variants), and multi-modal LLMs that use video-caption pretraining. LangRepo shows significantly-better performance compared to other methods at a similar scale, validating its effectiveness. We achieve $+7.8\%$ on fullset over mPLUG-Owl (Ye et al., 2023b), $+12.0\%$ on subset over pure Mistral LLM baseline (Jiang et al., 2023), $+10.0\%$ on subset and $+5.4\%$ on fullset over LLoVi (7B) (Zhang et al., 2023a) (w/ Mistral (Jiang et al., 2023)), $+4.5\%$ on fullset over Vamos (Wang et al., 2023) (w/ Llama2 (Touvron et al., 2023)), and $+4.8\%$ on subset over Tarsier (7B) (Wang et al., 2024a).
Table 3: Results on NExT-QA (Xiao et al., 2021): We compare LangRepo against state-of-the-art methods on NExT-QA validation set, highlighting standard splits: causal, temporal and descriptive. We focus on the zero-shot setting. Our method shows strong performance at its scale.
{tabu}
lcccccc  Model Video Pretrain Params  Causal (%)  Temporal (%)  Descriptive (%)  All (%)
finetuned
CoVGT (Xiao et al., 2023b) ✓ 149M 58.8 57.4 69.3 60.0
SeViTFiD (Kim et al., 2023) ✓ 215M - - - 60.6
HiTeA (Ye et al., 2023a) ✓ 297M 62.4 58.3 75.6 63.1
MC-ViT-L (Balažević et al., 2024) ✓ 424M - - - 65.0
InternVideo (Wang et al., 2022a) ✓ 478M 62.5 58.5 75.8 63.2
BLIP-2 (Li et al., 2023b) ✓ 4B 70.1 65.2 80.1 70.1
SeViLA (Yu et al., 2024) ✓ 4B 74.2 69.4 81.3 73.8
LLama-VQA (Ko et al., 2023) ✓ 7B 72.7 69.2 75.8 72.0
Vamos (Wang et al., 2023) ✓ 7B 72.6 69.6 78.0 72.5
zero-shot (with proprietary LLMs)
ViperGPT (Surís et al., 2023) ✗ 175B - - - 60.0
ProViQ (Choudhury et al., 2023) ✗ 175B - - - 64.6
MoReVQA (Min et al., 2024) ✗ 340B 70.2 64.6 - 69.2
LVNet (Park et al., 2024)$<$1.8T    75.0 65.5 81.5 72.9
IG-VLM (Kim et al., 2024) ✗ 1.8T 69.8 63.6 74.7 68.6
LLoVi (Zhang et al., 2023a) ✗ 1.8T 69.5 61.0 75.6 67.7
TraveLER (Shang et al., 2024) ✗ 1.8T 70.0 60.5 78.2 68.2
VideoAgent (Wang et al., 2024b) ✗ 1.8T 72.7 64.5 81.1 71.3
VideoTree (Wang et al., 2024e) ✗ 1.8T 75.2 67.0 81.3 73.5
zero-shot (with open-source LLMs)
VFC (Momeni et al., 2023) ✓ 164M 45.4 51.6 64.1 51.5
InternVideo (Wang et al., 2022a) ✓ 478M 43.4 48.0 65.1 49.1
SeViLA (Yu et al., 2024) ✓ 4B 61.3 61.5 75.6 63.6
Tarsier (Wang et al., 2024a) ✓ 7B - - - 71.6
Tarsier (Wang et al., 2024a) ✓ 34B - - - 79.2
Mistral (Jiang et al., 2023) ✗ 7B 51.0 48.1 57.4 51.1
LLoVi (Zhang et al., 2023a) ✗ 7B 55.6 47.9 63.2 54.3
LLoVi (Zhang et al., 2023a) ✗ 12B 60.2 51.2 66.0 58.2
LangRepo (ours) ✗ 7B 57.8 45.7 61.9 54.6
LangRepo (ours) ✗ 12B 64.4 51.4 69.1 60.9
@@ 5.1.0.2 NExT-QA:
In Table 5.1.0.1, we report the performance of LangRepo on standard NExT-QA (Xiao et al., 2021) validation splits (Causal, Temporal and Descriptive) and the full validation set. On zero-shot evaluation, our framework outperforms other methods consistently. Compared to smaller models, we gain $+11.8\%$ over InternVideo (Wang et al., 2022a) and $+9.4\%$ over VFC (Momeni et al., 2023). Compared to models of similar scale, we gain $+3.5\%$ over baseline Mistral LLM (Jiang et al., 2023) and $+2.7\%$ over LLoVi (12B) (Zhang et al., 2023a). We de-emphasize the comparisons with much-larger models, and multi-modal LLMs pretrained with video captions (whereas we rely on LLaVA-1.5 (Liu et al., 2023) captions that has not seen any video pretraining). Finally, we observe that LangRepo outperforms competition on semantic splits showing the generalization of our language representation.
Table 4: Results on IntentQA (Li et al., 2023a): We compare LangRepo against state-of-the-art methods on IntentQA test set, highlighting standard splits: why?, how? and before/after. We focus on the zero-shot setting. Our method shows strong performance at its scale.
{tabu}
lcccccc Model Video Pretrain Params  Why? (%)  How? (%)  Before/After (%)  All (%)
finetuned
HQGA (Xiao et al., 2022a) ✓ 46M 48.2 54.3 41.7 47.7
VGT (Xiao et al., 2022b) ✓ 511M 51.4 56.0 47.6 51.3
Vamos (Wang et al., 2023) ✓ 7B 69.5 70.2 65.0 68.5
BlindGPT (Ouyang et al., 2022) ✓ 175B 52.2 61.3 43.4 51.6
CaVIR (Li et al., 2023a) ✓ 175B 58.4 65.5 50.5 57.6
zero-shot (with proprietary LLMs)
LVNet (Park et al., 2024)$<$1.8T    75.0 74.4 62.1 71.7
LLoVi (Zhang et al., 2023a) ✗ 1.8T 68.4 67.4 51.1 64.0
IG-VLM (Kim et al., 2024) ✗ 1.8T - - - 64.2
VideoTree (Wang et al., 2024e) ✗ 1.8T - - - 66.9
zero-shot (with open-source LLMs)
SeViLA (Yu et al., 2024) ✓ 4B - - - 60.9
Mistral(Jiang et al., 2023) ✗ 7B 52.7 55.4 41.5 50.4
LLoVi (Zhang et al., 2023a) ✗ 7B 57.9 55.4 42.3 53.6
LLoVi (Zhang et al., 2023a) ✗ 12B 59.7 62.7 45.1 56.6
LangRepo (ours) ✗ 7B 56.9 60.2 42.1 53.8
LangRepo (ours) ✗ 12B 62.8 62.4 47.8 59.1